OpenStudy (anonymous):

Solve 7.74x + 3 = 10.2

OpenStudy (anonymous):

x ≈ –2.15 x ≈ –0.47 x ≈ 0.03 x ≈ 0.28

OpenStudy (anonymous):

\[7.7^{4x+8}=10.2\]

OpenStudy (anonymous):

@Teinis

OpenStudy (anonymous):

I got 0.93 but it isn't in the answer choices

OpenStudy (anonymous):

Ok I'm gunna try walk you through this one so you get the correct answer yourself.

OpenStudy (anonymous):

What did you do first?

OpenStudy (anonymous):

I really don't know I am confused :(

OpenStudy (anonymous):

Well ok. First off with take the log of both sides again

OpenStudy (anonymous):

(x+3)log[7.74] = log[10.2]. Divide both sides by log[7.74], giving x+3 = log[10.2]/log[7.74]. Subtract 3 from both sides, giving x = log[10.2]/log[7.74] - 3 ≈ -1.86

OpenStudy (anonymous):

\[\ln 7.7^{4x+8} = \ln10.2\]

OpenStudy (anonymous):

bUT -1.86 ISN'T AN ANSWER CHOICE

OpenStudy (anonymous):

see on the first step, you can't take out x+3 to the front like that

OpenStudy (anonymous):

you have to take out the whole thing, so (4x+3)

OpenStudy (anonymous):

Ok then is it this: Subtract 3 from both sides, giving 7.74^x = 7.2. Take logs, giving x·log[7.74] = log[7.2]. Divide both sides by log[7.74], giving x = log[7.2]/log[7.74] ≈ 0.96

OpenStudy (anonymous):

But 0.96 isn't an answer choice

OpenStudy (anonymous):

wait i'm confussed, is the equation \[7.7^{4x+3} = 10.2\] or \[7.7^4x + 3= 10.2\]

OpenStudy (anonymous):

Because if it's the first one then you cannot subtract 3 from both sides

OpenStudy (anonymous):

The first one

OpenStudy (anonymous):

Ok so then let's continue from after we take the logs

OpenStudy (anonymous):

move the whole 4x+3 to the front, so you get: \[(4x+3)\ln 7.7 = \ln10.2\]

OpenStudy (anonymous):

Divide both sides by ln 7.7

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

\[4x+3 = \frac{ \ln10.2 }{ \ln7.7}\]

OpenStudy (anonymous):

Now we can subtract 3 from both sides

OpenStudy (anonymous):

then we divide x?

OpenStudy (anonymous):

oh we subtract 3 and divide 4

OpenStudy (anonymous):

Yes exactly

OpenStudy (anonymous):

to get x alone i see

OpenStudy (anonymous):

so would the equation look like this

OpenStudy (anonymous):

Your first answer of -1.86 was almost correct actually, you just forgot to divide by 4 at the end

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

No problem